# STRIDE: Sparse Techniques for Regression in Deep Gaussian Processes

### v1

@@ 2.2 Deep Gaussian process regression
The posterior in deep GPR is commonly approximated with samples, e.g. through MCMC [14, 24, 34, 23]. Here we especially note a helpful trick in MCMC for deep GPR: marginalising the outer layer: Bayes’ theorem implies that
$\pi_{f_{L},\ldots,f_{0}|\mathbf{y}_{n}}=\frac{\pi_{\mathbf{y}_{n}|f_{L}}\pi_{f_{L},\ldots,f_{0}}}{\pi_{\mathbf{y}_{n}}}=\underbrace{\frac{\pi_{\mathbf{y}_{n}|f_{L}}\pi_{f_{L}|f_{L-1:0}}}{\pi_{\mathbf{y}_{n}|f_{L-1:0}}}}_{(*)}\underbrace{\frac{\pi_{\mathbf{y}_{n}|f_{L-1:0}}\pi_{f_{L-1:0}}}{\pi_{\mathbf{y}_{n}}}}_{(**)},$
where $f_{L-1:0}:=f_{L-1},\ldots,f_{0}$ and
    is the conditional outer layer posterior $\pi_{f_{L}|\mathbf{y}_{n},f_{L-1:0}}$ and given through (1),
    is the inner layers posterior $\pi_{f_{L-1:0}|\mathbf{y}_{n}}$ that can be sampled with MCMC.
Hence, the MCMC method needs to only approximate the posterior distributions on the lower layers $f_{L-1:0}$, but not the outer layer $f_{L}$.
@@ 2.3 Sparse Gaussian process regression
In GPR, the inversion of the kernel matrix becomes a computational challenge for large numbers of observations $n$. The idea of sparse GPR is to reduce the size of this matrix by representing the observed data through a number of “virtual” observations, or inducing variables. We denote the number of inducing variables by $m$, and generally choose $m\ll n$. We then construct an approximate posterior $\pi^{\mathcal{M}}_{f|\mathbf{y}_{n}}$ that only requires us to work with $n\times m$ and smaller matrices.
Suppose we have chosen a subset of the indices of the observational data points, $\mathcal{M}\subset[n]$ with $|\mathcal{M}|=m$, to represent our inducing points. Here, $[n]$ denotes the set $[n]=\{i\in\mathbb{N}\>|\>i\leq n\}$. Let $p:[m]\rightarrow[n]$ denote a mapping that assigns each observation selected as an inducing point its corresponding index in the data $(x_{i},y_{i})_{i=1}^{n}$. Using this mapping, define the matrices $K_{mm}\in\mathbb{R}^{m\times m}$ and $K_{mn}\in\mathbb{R}^{m\times n}$, as well as the vector of covariances $K_{m}(x)\in\mathbb{R}^{m}$, element-wise as
$(K_{mm})_{j,j^{\prime}}=k\mathopen{}\mathclose{{}\left(x_{p(j)},x_{p(j^{\prime})}}\right),\qquad(K_{mn})_{j,i}=k\mathopen{}\mathclose{{}\left(x_{p(j)},x_{i}}\right),\qquad(K_{m}(x))_{j}=k\mathopen{}\mathclose{{}\left(x_{p(j)},x}\right),$
where $i\in[n],j,j^{\prime}\in[m],x\in\mathbb{R}^{d}.$
In selecting the inducing points from the observations, we follow [38]. We choose inducing points represented by indices $\mathcal{M}$ that minimize the KL divergence between an approximate posterior $\pi^{\mathcal{M}}_{f|\mathbf{y}_{n}}$ and the posterior $\mathrm{GP}(\mu^{y}(\cdot),k^{y}(\cdot,\cdot))$. As is shown in [38, 37], this KL divergence can be minimised by maximising the variational lower bound,
$\begin{split}F_{V}(\mathcal{M})&\propto-\frac{1}{2}\mathbf{y}_{n}^{T}(K_{mn}^{T}K_{mm}^{-1}K_{mn}+\gamma^{2}\mathrm{Id}_{n})^{-1}\mathbf{y}_{n}-\frac{1}{2}\log\det(K_{mn}^{T}K_{mm}^{-1}K_{mn}+\gamma^{2}\mathrm{Id}_{n})\\ &\mathrel{\phantom{\propto}}-\frac{1}{2\gamma^{2}}\mathrm{trace}(K_{rr}-K_{mr}^{T}K_{mm}^{-1}K_{mr})\>.\end{split}$ (2)
@@
$K_{rr}=\mathopen{}\mathclose{{}\left(k\mathopen{}\mathclose{{}\left(x_{i},x_{j}}\right)}\right)_{i\in[n]\backslash\mathcal{M},\,j\in[n]\backslash\mathcal{M}},\qquad K_{mr}=\mathopen{}\mathclose{{}\left(k\mathopen{}\mathclose{{}\left(x_{i},x_{j}}\right)}\right)_{i\in\mathcal{M},j\in[n]\backslash\mathcal{M}}\>.$
The approximate posterior $\pi^{\mathcal{M}}_{f|\mathbf{y}_{n}}=\mathrm{GP}(\mu_{\mathcal{M}},k_{\mathcal{M}}(\cdot,\cdot))$ obtained with sparse GPR is given then by
$\mu_{\mathcal{M}}(x)$ $=\gamma^{-2}K_{m}(x)^{T}\mathopen{}\mathclose{{}\left(K_{mm}+\gamma^{-2}K_{mn}K_{mn}^{T}}\right)^{-1}K_{mn}\mathbf{y}_{n}\>,$
$k_{\mathcal{M}}(x,x^{\prime})$ $=k(x,x^{\prime})-K_{m}(x)^{T}K_{mm}^{-1}K_{m}(x^{\prime})+K_{m}(x)^{T}\mathopen{}\mathclose{{}\left(K_{mm}+\gamma^{-2}K_{mn}K_{mn}^{T}}\right)^{-1}K_{m}(x^{\prime})\>.$
@@ 3 Inducing points for deep Gaussian process regression
When using the marginalisation strategy in Subsection 2.2, we need to repeatedly evaluate $\pi_{\mathbf{y_{n}}|f_{L-1},\ldots,f_{0}}$. Each of the latter evaluations requires a GPR step, which is prohibitively expensive if $n$ is large. In the following, we introduce an iterative procedure that allows us find a set of inducing points $\mathcal{M}$ that help us to construct approximate distributions $\pi_{f_{L}|\mathbf{y_{n}},f_{L-1:0}}^{\mathcal{M}}$, $\pi_{f_{L-1:0}|\mathbf{y_{n}}}^{\mathcal{M}}$, and $\pi_{\mathbf{y_{n}}|f_{L-1:0}}^{\mathcal{M}}$ through sparse GPR and, thus, enable deep GPR under large-scale data.
@@
$F(\mathcal{M})=\int F_{V}(\mathcal{M};u_{L-1:0})\>\mathrm{d}\pi_{f_{L-1:0}|\mathbf{y_{n}}}^{\mathcal{M}}(u_{L-1:0}),$
where $F_{V}(\mathcal{M};u_{L-1:0})$ is given as in (2), but with the covariance kernel $k:=k_{L}$ depending on lower layers $u_{L-1:0}$. To minimise $F$, we propose a Monte Carlo Expectation-Maximisation (MCEM) method [39]. In MCEM methods, we maximises objectives such as $F$ by alternately approximating the distribution $\pi_{f_{L-1:0}|\mathbf{y_{n}}}^{\mathcal{M}}$ given some $\mathcal{M}$ (expectation step) and by using this approximation to maximise an empirical approximation of $F(\mathcal{M})$ and, thus, obtain an improved index set (maximisation step).
Let $P_{\mathcal{M}}$ be an MCMC kernel, i.e., a Markov kernel that is stationary with respect to $\pi_{f_{L-1:0}|\mathbf{y_{n}}}^{\mathcal{M}}$ and ergodic. We then construct a sequence of particles and inducing point sets $(f_{L-1:0}^{(1,t)},\ldots,f_{L-1:0}^{(S,t)},\mathcal{M}_{t})_{t\geq 0}$, where $S$ denotes the number of particles and $t$ denotes time. Particles and sets are given through the following two-step iteration.
  1. (E)
    Given inducing points $\mathcal{M}_{t}$, sample $f_{L-1:0}^{(s,t+1)}\sim P_{\mathcal{M}_{t}}\mathopen{}\mathclose{{}\left(\cdot\,|\,f_{L-1:0}^{(s,t)}}\right)$ for $s=1,...,S$.
  2. (M)
    Given particles $(f_{L-1:0}^{(s,t+1)})_{s=1}^{S}$, update the inducing points through
@@
We refer to the above method as STRIDE. Here, (E) corresponds to the expectation step, and the (M) to the maximisation step in this MCEM algorithm, respectively. We terminate the method at a large $T\in\mathbb{N}$, the final model is then given by
$\frac{1}{S}\sum_{s=1}^{S}\mathrm{GP}\mathopen{}\mathclose{{}\left(\mu_{\mathcal{M}_{T}}\mathopen{}\mathclose{{}\left(\cdot\,|\,f_{L-1:0}^{(s,T)}}\right),k_{\mathcal{M}_{T}}\mathopen{}\mathclose{{}\left(\cdot,\cdot\,|\,f_{L-1:0}^{(s,T)}}\right)}\right)\>.$
Both expectation and maximisation steps will be described in more detail in the following two subsections. Following this, we discuss computationally efficient strategies on the lower layers.
@@ 3.1 Monte Carlo expectation step
We first describe the procedure employed to obtain samples $(f_{L-1:0}^{(s,t+1)})_{s=1}^{S}$, as required in the expectation step (E) of STRIDE. Given some inducing points $\mathcal{M}_{t}$, we want sample the hidden parameter from the density $\pi_{f_{L-1:0}|\mathbf{y_{n}}}^{\mathcal{M}}$. This density is given through Bayes’ Theorem
$\pi_{f_{L-1:0}|\mathbf{y_{n}}}^{\mathcal{M}}\propto\pi_{\mathbf{y_{n}}|f_{L-1:0}}^{\mathcal{M}}\,\pi_{f_{L-1:0}}\ ,$
see also ()(**)( ∗ ∗ ). Combining the covariance structure defined by $f_{L-1:0}$ and $\mathcal{M}_{t}$ with the observational noise covariance, we can express the likelihood term as,
$\pi_{\mathbf{y_{n}}|f_{L-1:0}}^{\mathcal{M}}$ $\propto\frac{1}{\sqrt{\det Q_{nn}(f_{L-1:0})}}\>\exp\mathopen{}\mathclose{{}\left(-\frac{1}{2}\|\mathbf{y}_{n}\|^{2}_{Q_{nn}(f_{L-1:0})}}\right)\>,$
$Q_{nn}(f_{L-1:0})$ $=K_{mn}(f_{L-1:0})^{T}K_{mm}(f_{L-1:0})^{-1}K_{mn}(f_{L-1:0})+\gamma^{2}\mathrm{Id}_{n}\>.$
The logarithmic likelihood (omitting constants) is identical to the first two terms in our definition of $F_{V}$ in eq. 2. In order to compute samples of this density, we employ a Monte Carlo Markov chain (MCMC) algorithm. In particular, we whiten the prior $\pi_{f_{L-1:0}}$ [28, 43] and then apply the pCN proposal [6].
We note that the determinant of $Q_{nn}(f_{L-1:0})$ can be computed in $\mathcal{O}(m^{3})$ using the Weinstein–Aronszajn identity. We usually employ multiple subsequent MCMC steps in a single expectation step to ensure that the particles approximate $\pi_{f_{L-1:0}|\mathbf{y_{n}}}^{\mathcal{M}}$ well.

### v2

@@ 2.2 Deep Gaussian process regression
The posterior in deep GPR is commonly approximated with samples, e.g. through MCMC [dunlop18, monterrubio2020posterior, sauer2023active, latz2025deep]. Here we especially note a helpful trick in MCMC for deep GPR, which we refer to as marginalising the outer layer. Bayes’ theorem implies that
$\pi_{f_{L},\ldots,f_{0}|\mathbf{y}}=\frac{\pi_{\mathbf{y}|f_{L}}\pi_{f_{L},\ldots,f_{0}}}{\pi_{\mathbf{y}}}=\underbrace{\frac{\pi_{\mathbf{y}|f_{L}}\pi_{f_{L}|f_{L-1:0}}}{\pi_{\mathbf{y}|f_{L-1:0}}}}_{(*)}\underbrace{\frac{\pi_{\mathbf{y}|f_{L-1:0}}\pi_{f_{L-1:0}}}{\pi_{\mathbf{y}}}}_{(**)}\>,$
where $u_{L-1:0}:=u_{L-1},\ldots,u_{0}$ and
    is the conditional outer layer posterior $\pi_{f_{L}|\mathbf{y},f_{L-1:0}}$ and given through (1),
    is the inner layer’s posterior $\pi_{f_{L-1:0}|\mathbf{y}}$ that can be sampled with MCMC.
Hence, the MCMC method needs to approximate the posterior distributions on the lower layers $f_{L-1:0}$ only, but not the outer layer $f_{L}$.
We gain two advantages from marginalising the top layer. Firstly, one typically wants to avoid sampling where possible, as sampling tends to be computationally expensive. Crucially, the computational cost of most sampling methods increases with the dimension of the sampled quantity. However, there is another, more subtle benefit to treat the top layer separately. Typically, this top layer’s likelihood is particularly singular (concentrated) compared to the hidden layers. This means that sampling the top layer is more costly than sampling a hidden layer, as the associated Markov chain takes longer to reach stationarity. This is also what we observe in practice. If we sample the top layer, we need a prohibitively large number of samples to obtain a reasonable approximation to the top layer mean.
@@ 2.3 Sparse Gaussian process regression
In GPR, the inversion of the kernel matrix becomes a computational challenge for large numbers of observations $n$. Sparse GPR reduces the cost of inverting this matrix by representing the observed data through a much smaller number of virtual observations, or inducing variables. Here, the term virtual relates to the fact that generally, the inducing points do not have to be real observations, but can just be chosen to represent the observations well. However, in what follows, we only consider the case where the inducing points are selected from the observations. We denote the number of inducing variables by $m$, and generally choose $m\ll n$. We then construct an approximate posterior $\pi^{\mathcal{M}}_{f|\mathbf{y}}$ that only requires us to work with $n\times m$ and smaller matrices.
Suppose we have chosen a subset of the indices of the observational data points, $\mathcal{M}\subset[n]$ with $|\mathcal{M}|=m$, to represent our inducing points. Here, $[n]$ denotes the set $n]=\{i\in\mathbb{N}\>|\>i\leq n\}$. Let $\psi:[m]\rightarrow[n]$ denote a mapping that assigns each observation selected as an inducing point its corresponding index in the data $(\mathbf{x}_{i},y_{i})_{i=1}^{n}$. Using this mapping, define the matrices $K_{mm}\in\mathbb{R}^{m\times m}$ and $K_{mn}\in\mathbb{R}^{m\times n}$, as well as the vector of covariances $K_{m}(\mathbf{x})\in\mathbb{R}^{m}$, element-wise as
$(K_{mm})_{j,j^{\prime}}=k\mathopen{}\mathclose{{\left(\mathbf{x}_{\psi(j)},\mathbf{x}_{\psi(j^{\prime})}}}\right),\qquad(K_{mn})_{j,i}=k\mathopen{}\mathclose{{\left(\mathbf{x}_{\psi(j)},\mathbf{x}_{i}}}\right),\qquad(K_{m}(\mathbf{x}))_{j}=k\mathopen{}\mathclose{{\left(\mathbf{x}_{\psi(j)},\mathbf{x}}}\right),$
where $i\in[n],j,j^{\prime}\in[m],\mathbf{x}\in\mathbb{R}^{d}.$
In selecting the inducing points from the observations, we follow [titsias2009variational]. We choose inducing points represented by indices $\mathcal{M}$ that minimise the KL divergence between an approximate posterior $\pi^{\mathcal{M}}_{f|\mathbf{y}}$ and the posterior $\mathcal{GP}(\mu^{\mathbf{y}},k^{\mathbf{y}})$. As is shown in [titsias2009variational, titsias2009report], this KL divergence can be minimised by maximising the variational lower bound,
$\begin{split}F_{V}(\mathcal{M})&\propto-\frac{1}{2}\mathbf{y}^{T}(K_{mn}^{T}K_{mm}^{-1}K_{mn}+\gamma^{2}\mathrm{Id}_{n})^{-1}\mathbf{y}-\frac{1}{2}\log\det(K_{mn}^{T}K_{mm}^{-1}K_{mn}+\gamma^{2}\mathrm{Id}_{n})\\ &\mathrel{\phantom{\propto}}-\frac{1}{2\gamma^{2}}\mathrm{trace}(K_{rr}-K_{mr}^{T}K_{mm}^{-1}K_{mr})\>.\end{split}$ (3)
@@
$K_{rr}=\mathopen{}\mathclose{{\left(k\mathopen{}\mathclose{{\left(\mathbf{x}_{i},\mathbf{x}_{j}}}\right)}}\right)_{i\in[n]\backslash\mathcal{M},\,j\in[n]\backslash\mathcal{M}},\qquad K_{mr}=\mathopen{}\mathclose{{\left(k\mathopen{}\mathclose{{\left(\mathbf{x}_{i},\mathbf{x}_{j}}}\right)}}\right)_{i\in\mathcal{M},j\in[n]\backslash\mathcal{M}}\>$.
The approximate posterior $\pi^{\mathcal{M}}_{f|\mathbf{y}}=\mathcal{GP}(\mu_{\mathcal{M}},k_{\mathcal{M}})$ obtained with sparse GPR is then given by
$\mu_{\mathcal{M}}(\mathbf{x})$ $=\gamma^{-2}K_{m}(\mathbf{x})^{T}\mathopen{}\mathclose{{\left(K_{mm}+\gamma^{-2}K_{mn}K_{mn}^{T}}}\right)^{-1}K_{mn}\mathbf{y}\>,$
$k_{\mathcal{M}}(\mathbf{x},\mathbf{x}^{\prime})$ $=k(\mathbf{x},\mathbf{x}^{\prime})-K_{m}(\mathbf{x})^{T}K_{mm}^{-1}K_{m}(\mathbf{x}^{\prime})+K_{m}(\mathbf{x})^{T}\mathopen{}\mathclose{{\left(K_{mm}+\gamma^{-2}K_{mn}K_{mn}^{T}}}\right)^{-1}K_{m}(\mathbf{x}^{\prime})\>.$
@@ 3 Inducing points for deep Gaussian process regression
When using the marginalisation strategy in Subsection 2.2, we need to repeatedly evaluate $\pi_{\mathbf{y}|f_{L-1},\ldots,f_{0}}$. Each of the latter evaluations requires a GPR step, which is prohibitively expensive if $n$ is large. In the following, we introduce an iterative procedure that allows us find a set of inducing points $\mathcal{M}$ that help us to construct approximate distributions
$\pi_{f_{L}|\mathbf{y},f_{L-1:0}}^{\mathcal{M}},\quad\pi_{f_{L-1:0}|\mathbf{y}}^{\mathcal{M}},\quad\pi_{\mathbf{y}|f_{L-1:0}}^{\mathcal{M}}$
through sparse GPR and, thus, enable deep GPR under large-scale data.
@@
$F(\mathcal{M})=\int F_{V}(\mathcal{M};u_{L-1:0})\>\mathrm{d}\pi_{f_{L-1:0}|\mathbf{y}}^{\mathcal{M}}(u_{L-1:0})\>,$ (4)
where $F_{V}(\mathcal{M};u_{L-1:0})$ is given as in (3), but with the covariance kernel $k:=k_{L}$ depending on lower layers $u_{L-1:0}$. To maximise $F$, we propose a Monte Carlo Expectation-Maximisation (MCEM) method, see [wei1990monte] for details.
Expectation-Maximisation (EM) methods are used to maximise an objective that is typically parametrised by missing data. In an expectation step, we infer the missing data given an estimate of the optimum. In a maximisation step, this inferred missing data is then used to approximate the objective and find the next estimate of the optimum through optimisation. Both steps are re-iterated alternately until a convergence criterion is satisfied. In our case, the objective $F$ depends on the distribution represented via the density $\pi_{f_{L-1:0}|\mathbf{y}}^{\mathcal{M}}$. Thus, in the expectation step, we aim at estimating the expected value as given in eq.˜4 for a given set of inducing points $\mathcal{M}$. In the maximisation step, we optimise the approximate objective obtained in the expectation step. However, due to the nature of our problem, we are not able to evaluate $F(\mathcal{M})$ exactly. In particular, we rely on MCMC to sample from $\pi_{f_{L-1:0}|\mathbf{y}}^{\mathcal{M}}$ and obtain an empirical approximation of $F(\mathcal{M})$ in the expectation step, making our method an MCEM algorithm.
Let $P_{\mathcal{M}}$ be an MCMC kernel, i.e., a Markov kernel that is stationary with respect to $\pi_{f_{L-1:0}|\mathbf{y}}^{\mathcal{M}}$ and ergodic. We then construct a sequence of particles and inducing point sets $\mathopen{}\mathclose{{\left(f_{L-1:0}^{(1,t)},\ldots,f_{L-1:0}^{(S,t)},\mathcal{M}_{t}}}\right)_{t\geq 0}$, where $S$ denotes the number of particles and $t$ denotes time. Particles and sets are given through the following two-step iteration.
  • (E)
    Given inducing points $\mathcal{M}_{t}$, sample
    $f_{L-1:0}^{(s,t+1)}\sim P_{\mathcal{M}_{t}}\mathopen{}\mathclose{{\left(\cdot\,\Big|\,f_{L-1:0}^{(s,t)}}}\right)\qquad(s\in[S]).$
  • (M)
    Given particles $\mathopen{}\mathclose{{\left(f_{L-1:0}^{(s,t+1)}}}\right)_{s=1}^{S}$, update the inducing points through
@@
We refer to the above method as STRIDE. Here, (E) corresponds to the expectation step, and the (M) to the maximisation step in this MCEM algorithm, respectively. We terminate the method at a large $T\in\mathbb{N}$. The final model is then given by
$\frac{1}{S}\sum_{s=1}^{S}\mathcal{GP}\mathopen{}\mathclose{{\left(\mu_{\mathcal{M}_{T}}\mathopen{}\mathclose{{\left(\cdot\,\Big|\,f_{L-1:0}^{(s,T)}}}\right),k_{\mathcal{M}_{T}}\mathopen{}\mathclose{{\left(\cdot,\cdot\,\Big|\,f_{L-1:0}^{(s,T)}}}\right)}}\right)\>.$ (5)
We summarise our algorithm in Algorithm˜1. Both the expectation and the maximisation step will be described in more detail in the following two subsections. Following this, we discuss computationally efficient strategies on the lower layers.
Input: Number of particles $S$, maximum time value $T$, initial hidden parameters $f_{L-1:0}^{(s,0)}$, observations $(X,\mathbf{y})$, noise level $\gamma^{2}$.
Select initial inducing points $\mathcal{M}_{0}\in\mathrm{argmax}_{\mathcal{M}}\frac{1}{S}\sum_{s=1}^{S}F_{V}\mathopen{}\mathclose{{\left(\mathcal{M};f_{L-1:0}^{(s,0)}}}\right)$.
for $t=0,~1,~\ldots,~T-1$ do
 (E)(M)Sample $f_{L-1:0}^{(s,t+1)}\sim P_{\mathcal{M}_{t}}\mathopen{}\mathclose{{\left(\cdot\,\Big|\,f_{L-1:0}^{(s,t)}}}\right)\quad(s\in[S])$.
 (M)(E)Update inducing points to obtain $\mathcal{M}_{t+1}$.
Output: Samples of hidden parameters $f_{L-1:0}^{(s,T)}$, inducing points $\mathcal{M}_{T}$.
Algorithm 1 STRIDE.
@@ 3.1 Monte Carlo expectation step
We first describe the procedure employed to obtain samples $\mathopen{}\mathclose{{\left(f_{L-1:0}^{(s,t+1)}}}\right)_{s=1}^{S}$, as required in the expectation step (E) of STRIDE. Given some inducing points $\mathcal{M}_{t}$, we want to sample the hidden parameter from the density $\pi_{f_{L-1:0}|\mathbf{y}}^{\mathcal{M}_{t}}$. This density is given through Bayes’ Theorem
$\pi_{f_{L-1:0}|\mathbf{y}}^{\mathcal{M}_{t}}\propto\pi_{\mathbf{y}|f_{L-1:0}}^{\mathcal{M}_{t}}\,\pi_{f_{L-1:0}}\>,$
see also ()(**). Combining the covariance structure defined by $f_{L-1:0}$ and $\mathcal{M}_{t}$ with the observational noise covariance, we can express the likelihood term as,
$\pi_{\mathbf{y}|f_{L-1:0}}^{\mathcal{M}_{t}}$ $\propto\frac{1}{\sqrt{\det Q_{nn}^{L-1:0}}}\>\exp\mathopen{}\mathclose{{\left(-\frac{1}{2}\|\mathbf{y}\|^{2}_{Q_{nn}^{L-1:0}}}}\right)\>,$
where
$\mathrel{\phantom{=}}Q_{nn}^{L-1:0}$ $=K_{mn}(f_{L-1:0})^{T}K_{mm}(f_{L-1:0})^{-1}K_{mn}(f_{L-1:0})+\gamma^{2}\mathrm{Id}_{n}\>.$
The logarithmic likelihood (omitting constants) is identical to the first two terms in our definition of $F_{V}$ in eq.˜3. In order to compute samples of this density, we employ an MCMC algorithm as described in, e.g., [cotter2013mcmc, dunlop18, latz2025deep, yu2011center]. In particular, we whiten the prior given by the density $\pi_{f_{L-1:0}}$ [papaspiliopoulos2007general, yu2011center] and then apply the preconditioned Crank–Nicolson (pCN) proposal [cotter2013mcmc].
We note that the determinant of $Q_{nn}^{L-1:0}$ can be computed in $\mathcal{O}(m^{3})$ using the Weinstein–Aronszajn identity. We usually employ multiple subsequent MCMC steps in a single expectation step to ensure that the particles approximate $\pi_{f_{L-1:0}|\mathbf{y}}^{\mathcal{M}_{t}}$ well.